OpenStudy (anonymous):

You invest $6000 into an account paying 3.75% annual interest compounded bimonthly(every 2 months), and you want to have $10,000 saved for your first year of college. How long will it take for your investment to grow that large? Round your answer to the nearest whole number. Answer a. 14 b. 15 c. 13 d. 12

OpenStudy (kropot72):

\[6000\times (1+\frac{0.0375}{6})^{x}=10000\] \[(1.00625)^{x}=\frac{10000}{6000}\]\[x \times \ln 1.00625=\ln \frac{10}{6}\] \[x=\frac{\ln \frac{10}{6}}{\ln 1.00625}=82\] Number of years needed = 82/6 = 14 years (rounded)